A 2.96 L sample of N2 gas at 271.4 K is heated to 327.6 K. What is the new volume (in L) of the N2 sample? Enter only the numeric value for your answer (no units).
Answer #1

Given:

Vi = 2.96 L

Ti = 271.4 K

Tf = 327.6 K

use:

Vi/Ti = Vf/Tf

2.96 L / 271.4 K = Vf / 327.6 K

Vf = 3.57 L

Answer: 3.57